OpenStudy (anonymous):

4/9 divided by 2/3 divided by 5/6

OpenStudy (anonymous):

4/9 * 3/2 = 2/3 2/3 * 6/5 = 12/15 = 4/5

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

divide by 2/3 = multiply by 3/2 - invert the divisor and multiply